# HumanSplat: Generalizable Single-Image Human Gaussian Splatting with Structure Priors

### v1

@@ 3.3 Video Diffusion Model as Novel-view Synthesizer
We take advantage of the pre-trained video diffusion model SV3D voleti2024sv3d as appearance prior. It utilizes the CLIP CLIP:2021 image embedding $\mathbf{c}$ and latent feature $\mathbf{F}_{0}$ of a input image $\mathbf{I}_{0}$ from a pre-trained VQ-VAE VQGAN:2023 $\mathcal{E}$ as conditions and progressively denoises gaussian noises into temporal-continuous $\mathbf{N}$ multi-view latent features $\{\mathbf{F}_{i}\}_{i=1}^{\mathbf{N}}$ by a spatial-temporal UNet $D_{\theta}$ SVD:2023 with the objective:
$\mathbb{E}_{\epsilon\sim p(\epsilon)}\left[\lambda(\epsilon)\|D_{\theta}(\{\mathbf{F}_{i}^{\epsilon}\}_{i=1}^{\mathbf{N}};\mathbf{c},\mathbf{F}_{0},\epsilon)-\{\mathbf{F}_{i}\}_{i=1}^{\mathbf{N}}\|^{2}_{2}\right],$ (1)
where $\{\mathbf{F}_{i}^{\epsilon}\}$ is the noise-corrupt multi-view latent features, $p(\epsilon)$ is the noise probability distribution and $\lambda(\epsilon)\in\mathbb{R}^{+}$ is the loss weighting term related to the noise level $\epsilon$. Since SV3D is merely pre-trained on the dataset of objects deitke2023objaverse , we further fine-tune it using human datasets to improve its effectiveness in the human reconstruction task. More technical details are available in Appendix A.1.
3.4 Latent Reconstruction Transfomer
Latent reconstruction transformer contains two parts: latent embedding interaction (Sec. 3.4.1) and geometry-aware interaction (Sec. 3.4.2), which seamlessly incorporate latent features from video diffusion and human structure priors into the reconstruction process. The two components together form a cohesive framework for accurately recovering intricate human details.
@@ 3.4.1 Latent Embedding Interaction
We utilize a pretrained VQ-VAE VQGAN:2023 $\mathcal{E}$ to encode the input image $\mathbf{I_{0}}$ into a latent representation $\mathbf{F}_{0}=\mathcal{E}(\mathbf{I_{0}})\in\mathbb{R}^{h\times w\times c}$ and combine it with the generated $\left\{\mathbf{F}_{i}\right\}_{i=1}^{\mathbf{N}}$ (Sec. 3.3). Following previous works xu2023dmv3d ; chen2023:ray-conditioning , Plücker ray embeddings plucker1865xvii are leveraged for pose conditioning, where latent representations are concatenated with their corresponding Plücker coordinates along the channel dimension, resulting in a per-pixel pose-conditioned feature map. These latents are then divided into non-overlapping patches dosovitskiy2020image and mapped to $d$-dimensional latent tokens by a linear layer. An intra-attention module is followed to thoroughly extract spatial correlations of latent tokens, as shown in Fig.3. The intra-attention module is implemented using a standard Transformer vaswani2017attention with multiple blocks of multi-head self-attention and feed-forward network (FFN):
$\bar{\mathbf{F}}_{0},\bar{\mathbf{F}}_{1},...,\bar{\mathbf{F}}_{\mathbf{N}}=\text{FFN}(\text{SelfAttention}(\mathbf{F}_{0},\mathbf{F}_{1},...,\mathbf{F}_{\mathbf{N}})).$ (2)
@@ 3.4.2 Geometry-aware Interaction
Human Geometric Tokenizer. Given a human geometric model $\mathcal{M}\in\mathbb{R}^{6890\times 3}$, to obtain the feature representation at a position $x\in\mathbb{R}^{3}$, we project $x$ onto input view and computes the feature vector through bilinear interpolation on the feature grids by
$\mathbf{\Pi}_{i}(x)=\mathbf{K}\left(\mathbf{R}\mathcal{M}_{i}+\mathbf{t}\right),$ (3)
@@
where $\mathbf{R,t}$ are camera extrinsic parameters, $\mathbf{K}$ is the intrinsic parameters. After the projection operation, we obtain geometric human prior tokens by concatenating $x$ and $\mathbf{F}_{0}(\mathbf{\Pi}(x))$, and mapping them to $d$ dimension. Similar to Latent Embedding Interaction, an intra-attention module is utilized to interact among these human prior tokens and get geometric features $\bar{\mathbf{H}}_{i}\in\mathbb{R}^{6890\times d}$.
@@
Human Geometry-aware Attention. Empirically, we find that incorporating human priors SMPL:2015 ; SMPL-X:2019 ; SMPLD:2019 (e.g., SMPL model) into human reconstruction faces a significant dilemma of the trade-off between robustness and flexibility GTA:2024 ; SIFU:2024 . On the one hand, the SMPL model provides geometric priors that alleviate common issues (e.g., broken body parts and various artifacts). On the other hand, it struggles to accurately capture and represent a broad spectrum of complex clothing styles, particularly more fluid garments like dresses and skirts. This limitation highlights a fundamental shortcoming in the model’s capacity to accurately depict diverse apparel. As shown in Fig. 3, HumanSplat efficiently utilizes priors while accounting for their inaccuracies by projecting 3D tokens into 2D space and searching in the nearby windows. To tackle the issue of matching redundancy with erroneous human priors, HumanSplat leverages these priors while compensating for their inaccuracies by projecting 3D tokens into 2D space and conducting searches within adjacent windows. Specifically, we introduce project-aware attention within the inter-attention module, utilizing a window $\mathbf{W}(k_{win}\times k_{win})$ and employing masked multi-head cross-attention. Here, the latent features $\left\{\bar{\mathbf{F}}_{i}\right\}$ serve as queries, and the human prior tokens $\left\{\bar{\mathbf{H}}_{i}\right\}$ serve as keys/values. The process can be formulated as follows:
$\tilde{\mathbf{F}}_{i}=\text{FFN}(\text{CrossAttention}_{mask}(\bar{\mathbf{F}}_{i},\bar{\mathbf{H}}_{i})),$ (4)
@@
where feature interactions occur when $\mathbf{\Pi}(\bar{\mathbf{H}}_{i},\mathbf{K},\mathbf{R},\mathbf{t})$ is within the window of $\bar{\mathbf{F}}_{i}$. It is noteworthy that the complexity of the cross-attention operation is reduced from $\mathcal{O}(\mathbf{L}_{F}\times\mathbf{L}_{H})$ to $\mathcal{O}(\mathbf{L}_{F}\times k_{win}^{2})$, compared to the vanilla multi-head Cross-attention setting. Here, $\mathbf{L}_{F}$ and $\mathbf{L}_{H}$ represent the lengths of the latent and human prior tokens respectively.
@@
Figure 3: Illustration of latent reconstruction transformer. It first divides $\mathbf{F}_{0}$ and $\mathbf{F}_{i}$ into non-overlapping patches, which are then processed through an intra-attention module (Sec. 3.4.1). Within the iter-attention module (Sec. 3.4.2), we introduce the project-aware attention with a window $\mathbf{W}(k_{win}\times k_{win})$, and the attributes of 3D Gaussians are decoded with a Conv $1\times 1$ layer.
@@ 3.5 Semantics-guided Objectives
From each output token $\tilde{\mathbf{F}}_{i}$, we decode the attributes of pixel-aligned Gaussians $\mathbf{\mathbf{G}}$ in the corresponding patch by a convolutional layer with a $1\times 1$ kernel. These attributes are used to render novel view images through 3D Gaussian Splatting. An ideal training objective should ensure that the rendered outputs closely match the supervised images, striving for overall consistency between the reconstructed renderings and the ground truth. However, human attention tends to be particularly focused on facial regions, where intricate details play a pivotal role in perception. Therefore, in our quest to elevate the fidelity of reconstructed human models, we introduce a paradigm shift in our training objectives.
@@
Hierarchical Loss. Traditional approaches often overlook the semantic richness inherent in human anatomy, leading to reconstructions that lack crucial details and accuracy. To address this, we propose a novel framework that harnesses semantic cues, hierarchical supervision, and tailored loss functions to guide the training process. This ensures that not only are the overall structures of the human body accurately represented, but also that critical areas such as the face are reconstructed with exceptional detail and fidelity. Specifically, by incorporating a human prior model equipped with semantic information, we establish direct correspondences between 3D spatial positions and specific human body parts. Therefore, we can utilize different attention weights for different body parts and render different levels of GT images from different viewpoints to provide hierarchical supervision. This approach significantly facilitates the precise localization of essential body parts, including the head, hands, and arms. The overall loss is defined as a weighted sum of part-specific losses:
$\mathcal{L_{H}}=\frac{1}{l}\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{m}\lambda_{i}\lambda_{j}\mathcal{L}_{Rec}(\mathbf{I}_{\text{part},j},\hat{\mathbf{I}}_{\text{part},j}),$ (5)
@@
where $i\in\left\{1,...,n\right\}$ and $j\in\left\{1,...,m\right\}$ denote different resolution levels and human parts. $\hat{\mathbf{I}}_{\text{part},j}$ and $\mathbf{I}_{\text{part},j}$ are the predicted and ground-truth data for part $j$, $\lambda_{i}$ and $\lambda_{j}$ are weighting factors that reflect the relative importance of each resolution level and each body part, respectively.
@@
Reconstruction Loss. Reconstruction loss $L_{Rec}$ based on the $\mathbf{N}_{\text{render}}$ rendered multi-view images is defined as
$\mathcal{L}_{Rec}=\sum_{i}^{\mathbf{N}_{\text{render}}}\mathcal{L}_{mse}(\hat{\mathbf{I}}_{i},\mathbf{I}_{i})+\lambda_{m}L_{mse}(\hat{\mathbf{M}}_{i},\mathbf{M})+\lambda_{p}\mathcal{L}_{p}(\hat{\mathbf{I}}_{i},\mathbf{I}_{i}),$ (6)
@@
where $\mathbf{I}_{i}$ and $\hat{\mathbf{I}}_{i}$ are the ground-truth images and render images via Gaussian Splatting, $\mathbf{M}_{i}$ and $\hat{\mathbf{M}}_{i}$ are original and rendered foreground mask. $\mathcal{L}_{mse}$ measures the MSE loss and $\mathcal{L}_{p}$ measures the perceptual loss (LIPIS:2018, ). $\lambda_{m}$ and $\lambda_{p}$ are hyperparameters for balancing. The final training objective is a combination of the two loss terms: $\mathcal{L}=\mathcal{L}_{H}+\mathcal{L}_{Rec}$.
@@ 4.1 Implementation Details
4.2 Comparison
Baselines. We evaluate novel view synthesis using state-of-the-art methods such as GTA GTA:2024 and SIFU SIFU:2024 by rendering textured meshes from multiple angles and calculating the PSNR to ground truth images. We divided 21 scans from Thuman2.0 THuman2.0:2021 and 51 scans from Twindom Twindom for evaluation. To eliminate the bias of the training dataset, we fine-tune the generative model LGM LGM:2024 on the same datasets as ours for a fair comparison. Additionally, we optimized each instance of TeCH TECH:2023 until full convergence for comparison.
[…]
Qualitative Comparison. By incorporating semantics-guided objectives, Humansplat achieves more detailed and higher fidelity results against GTA GTA:2024 and LGM LGM:2024 , as illustrated in Fig. 4. Results on in-the-wild images demonstrate our strong generalization capability. Although TeCH TECH:2023 also generates high quality image, its SDS optimization often results in overly saturated multi-face outcomes poole2022dreamfusion ; wang2023score . Moreover, while TeCH delivers highest PSNR on the Thuman2.0 dataset, the LPIPS metric in Tab. 1, confirms that our method delivers more realistic results. Notably, our model provides superior textures on both the front and invisible regions, highlighting the effectiveness of incorporating appearance and geometric priors.

### v2

@@ 3.3 Video Diffusion Model as Novel-view Synthesizer
We take advantage of the pre-trained video diffusion model SV3D [93] as appearance prior. For the input image $\mathbf{I}_{0}$, we leverage a CLIP image encoder [63] to obtain the image embedding $\mathbf{c}$ and a pre-trained VAE [94] $\mathcal{E}$ to acquire latent feature $\mathbf{F}_{0}$ as conditions. Subsequently, we progressively denoises gaussian noises into temporal-continuous $\mathbf{N}$ multi-view latent features $\{\mathbf{F}_{i}\}_{i=1}^{\mathbf{N}}$ by a spatial-temporal UNet $D_{\theta}$ [95] with the objective:
$\mathbb{E}_{\epsilon\sim p(\epsilon)}\left[\lambda(\epsilon)\|D_{\theta}(\{\mathbf{F}_{i}^{\epsilon}\}_{i=1}^{\mathbf{N}};\{e_{i},a_{i}\}_{i=1}^{\mathbf{N}},\mathbf{c},\mathbf{F}_{0},\epsilon)-\{\mathbf{F}_{i}\}_{i=1}^{\mathbf{N}}\|^{2}_{2}\right],$ (1)
where $\{\mathbf{F}_{i}^{\epsilon}\}_{i=1}^{\mathbf{N}}$ is the noise-corrupted multi-view latent features, $\{e_{i},a_{i}\}_{i=1}^{\mathbf{N}}$ is the corresponding relative elevation and azimuth angles to $\mathbf{I}_{0}$, $p(\epsilon)$ is the noise probability distribution and $\lambda(\epsilon)\in\mathbb{R}^{+}$ is the loss weighting term related to the noise level $\epsilon$. Since SV3D is merely pre-trained on the dataset of objects [96], we further fine-tune it using human datasets to improve its effectiveness in the human reconstruction task. More technical details are available in Appendix A.1.
3.4 Latent Reconstruction Transformer
Latent reconstruction Transformer contains two parts: latent embedding interaction (Sec. 3.4.1) and geometry-aware interaction (Sec. 3.4.2), which seamlessly incorporate latent features from novel-view synthesizer and human structure priors into the reconstruction process. The two components together form a cohesive framework for accurately recovering intricate human details.
@@ 3.4.1 Latent Embedding Interaction
We utilize a pretrained VAE [94] encoder $\mathcal{E}$ to encode the input image $\mathbf{I_{0}}$ into a latent representation $\mathbf{F}_{0}=\mathcal{E}(\mathbf{I_{0}})\in\mathbb{R}^{h\times w\times c}$ and combine it with the generated $\left\{\mathbf{F}_{i}\right\}_{i=1}^{\mathbf{N}}$ (Sec. 3.3). Following previous works [97, 98], latent representations are concatenated with their corresponding Plücker embeddings [99] along the channel dimension, resulting in a dense pose-conditioned feature map. These latents are then divided into non-overlapping patches [100] and mapped to $d$-dimensional latent tokens by a linear layer. An intra-attention module is followed to thoroughly extract spatial correlations of latent tokens, as shown in Fig. 3, by a standard Transformer [76] with $N_{\text{intra}}$ blocks of multi-head self-attention and feed-forward network (FFN):
$\bar{\mathbf{F}}_{i}=[\text{FFN}(\text{SelfAttention}(\mathbf{F}_{i}))]_{\times N_{\text{intra}}}.$ (2)
@@ 3.4.2 Geometry-aware Interaction
Human Geometric Tokenizer. Given a human geometric model $\mathcal{M}\in\mathbb{R}^{6890\times 3}$, to obtain the feature representation, we project $\mathcal{M}$ onto the input view and compute the feature vector through bilinear interpolation on the feature grids by
$\mathbf{\Pi}_{0}(\mathcal{M})=\mathbf{K}\left(\mathbf{R}_{0}\mathcal{M}+\mathbf{t}_{0}\right),$ (3)
@@
where $\mathbf{R}_{0}$, $\mathbf{t}_{0}$ are camera extrinsic parameters of the input view, $\mathbf{K}$ is the intrinsic parameters. After the projection operation, we obtain geometric human prior tokens by concatenating $\mathcal{M}$ and $\mathbf{F}_{0}[\mathbf{\Pi}_{0}(\mathcal{M})]$, and mapping them to $d$ dimension. $\mathbf{F}_{0}[\mathbf{\Pi}_{0}(\mathcal{M})]$ stands for querying the input latent feature $\mathbf{F}_{0}$ from projected locations. Similar to Latent Embedding Interaction (Sec. 3.4.1), an intra-attention module is utilized to interact among these human prior tokens and get human geometric features $\bar{\mathbf{H}}\in\mathbb{R}^{6890\times d}$.
@@
Human Geometry-aware Attention. Empirically, we find that incorporating human priors [5, 6, 101] (e.g., SMPL model) into human reconstruction faces a significant dilemma of the trade-off between robustness and flexibility [30, 102]. On the one hand, the SMPL model provides geometric priors that alleviate common issues (e.g., broken body parts and various artifacts). On the other hand, it struggles to accurately capture and represent a broad spectrum of complex clothing styles, particularly more fluid garments like dresses and skirts. This limitation highlights a fundamental shortcoming in the model’s capacity to accurately depict diverse apparel. As shown in Fig. 3, HumanSplat addresses inaccuracies in human priors by projecting 3D tokens into 2D space and conducting local searches within adjacent windows, efficiently using priors while minimizing redundancy. Specifically, we introduce projection-aware attention within the inter-attention module, utilizing a window $\mathbf{W}(K_{\text{win}}\times K_{\text{win}})$ and employing masked multi-head cross-attention. Here, the latent features $\left\{\bar{\mathbf{F}}_{i}\right\}_{i=1}^{\mathbf{N}}$ serve as queries, and the human prior tokens $\bar{\mathbf{H}}$ serve as keys and values. The process can be formulated as follows:
$\{\tilde{\mathbf{F}}_{i}\}_{i=1}^{\mathbf{N}}=[\text{FFN}(\text{CrossAttention}_{\text{mask}}(\{\bar{\mathbf{F}}_{i}\}_{i=1}^{\mathbf{N}},\bar{\mathbf{H}}))]_{\times N_{\text{inter}}},$ (4)
@@
where feature interactions occur when $\mathbf{\Pi}(\bar{\mathbf{H}},\mathbf{K},\mathbf{R}_{i},\mathbf{t}_{i})$ is within the window of $\bar{\mathbf{F}}_{i}$. It is noteworthy that the complexity of the cross-attention operation is reduced from $\mathcal{O}(\mathbf{L}_{F}\times\mathbf{L}_{H})$ to $\mathcal{O}(\mathbf{L}_{F}\times K_{\text{win}}^{2})$, compared to the vanilla multi-head cross-attention setting, where $\mathbf{L}_{F}$ and $\mathbf{L}_{H}$ represent the lengths of the latent and human geometric tokens respectively.
@@
Figure 3: Illustration of latent reconstruction Transformer. It first divides $\mathbf{F}_{0}$ and $\mathbf{F}_{i}$ into non-overlapping patches, which are then processed through an intra-attention module (Sec. 3.4.1). Within the iter-attention module (Sec. 3.4.2), we introduce the projection-aware attention with a window $\mathbf{W}(K_{\text{win}}\times K_{\text{win}})$, and the attributes of 3D Gaussians are decoded with a Conv $1\times 1$ layer.
@@ 3.5 Semantics-guided Objectives
From each output token $\tilde{\mathbf{F}}_{i}$, we decode the attributes of pixel-aligned Gaussians $\mathbf{\mathbf{G}}$ in the corresponding patch by a convolutional layer with a $1\times 1$ kernel. These attributes are used to render novel-view images through 3D Gaussian Splatting rasterization. An ideal training objective should ensure that the rendered outputs closely match the supervised images, striving for overall consistency between the reconstructed renderings and the ground truth. However, human attention tends to be particularly focused on facial regions, where intricate details play a pivotal role in perception. Therefore, to enhance the fidelity of reconstructed human models, we optimize our objective functions to preferentially focus on the facial regions.
@@
Hierarchical Loss. Traditional approaches often overlook the semantic richness inherent in human anatomy, leading to reconstructions that lack crucial details and accuracy. To address this, we propose a novel framework that harnesses semantic cues, hierarchical supervision, and tailored loss functions to guide the training process. This ensures that not only are the overall structures of the human body accurately represented, but also that critical areas such as the face are reconstructed with exceptional detail and fidelity. Specifically, by incorporating a human prior model equipped with semantic information, which adheres to the widely-used definition of 24 human body parts, such as “right hand”, “left hand”, “head”, etc. The ground-truth body part segmentation is obtained from SMPL meshes with predefined semantic vertices [103]. Therefore, we can utilize different attention weights for different body parts and render different levels of real images from different viewpoints to provide hierarchical supervision. This approach significantly facilitates the precise localization of essential body parts, including the head, hands and arms. Please refer to Appendix A.2 for more detailed procedural insights. The overall loss is defined as a weighted sum of part-specific losses:
$\mathcal{L}_{H}=\frac{1}{n}\frac{1}{m}\sum_{i=1}^{n}\sum_{j=1}^{m}\lambda_{i}\lambda_{j}(\mathcal{L}_{\text{MSE}}(\mathbf{I}_{i,j},\hat{\mathbf{I}}_{i,j})+\lambda_{p}\mathcal{L}_{\text{p}}(\mathbf{I}_{i,j},\hat{\mathbf{I}}_{i,j})),$ (5)
@@
where $i\in\left\{1,...,n\right\}$ and $j\in\left\{1,...,m\right\}$ denote different resolution levels and human parts. $\hat{\mathbf{I}}_{\text{i},j}$ and $\mathbf{I}_{i,j}$ are the predicted and ground-truth data for part $j$, $\lambda_{i}$ and $\lambda_{j}$ are weighting factors that reflect the relative importance of each resolution level and each body part, respectively. $\mathcal{L}_{\text{MSE}}$ measures the MSE loss and $\mathcal{L}_{p}$ measures the perceptual loss [104].
@@
Reconstruction Loss. Reconstruction loss $\mathcal{L}_{\text{{Rec}}}$ based on the $\mathbf{N}_{\text{render}}$ rendered multi-view images is defined as
$\mathcal{L}_{\text{Rec}}=\sum_{i=1}^{\mathbf{N}_{\text{render}}}\mathcal{L}_{\text{{MSE}}}(\mathbf{I}_{i},\hat{\mathbf{I}}_{i})+\lambda_{m}\mathcal{L}_{\text{MSE}}(\mathbf{M},\hat{\mathbf{M}}_{i})+\lambda_{p}\mathcal{L}_{\text{p}}(\mathbf{I}_{i},\hat{\mathbf{I}}_{i}),$ (6)
@@
where $\mathbf{I}_{i}$ and $\hat{\mathbf{I}}_{i}$ are the ground-truth images and render images via Gaussian Splatting, $\mathbf{M}_{i}$ and $\hat{\mathbf{M}}_{i}$ are original and rendered foreground mask. $\lambda_{m}$ and $\lambda_{\text{p}}$ are hyperparameters for balancing three kinds of losses. The final training objective is a combination of the two loss terms: $\mathcal{L}=\mathcal{L}_{H}+\mathcal{L}_{\text{Rec}}$.
[…]
Method Category THuman2.0 [105] Twindom [106] Time
[…]
PIFu [107] 18.093 0.911 0.137 - - - 30.0s
LGM [29] 20.013 0.893 0.116 19.840 0.851 0.292 9.5s
GTA [30] 18.050 - - 17.669 0.741 0.418 43s
SIFU [102] 22.025 0.921 0.084 19.714 0.832 0.312 44s
SIFU [102] 22.102 0.923 0.079 - - - 6min
Magic123 [108] 14.501 0.874 0.145 - - - 1h
HumanSGD [22] 17.365 0.895 0.130 - - - 7min
TeCH [21] 25.211 0.936 0.083 21.192 0.884 0.188 4.5h
HumanSplat (Ours) 24.033 0.918 0.055 23.346 0.913 0.125 9.3s
Figure 4: Qualitative comparison of ours against TeCH [21], GTA [30] and LGM [29] on Thuman2.0 [105], Twindom [106] and in-the-wild images. Our method achieves the highest quality. Note that TeCH achieves clearer results but fails to preserve the face identity.
@@ 4.1 Implementation Details
Inference Time. The video diffusion model takes about 9s to generate multi-view latent features while the subsequent latent reconstruction for 3DGS takes only about 0.3s. Additionally, it can render novel views at a rate exceeding 150 FPS on a NVIDIA A100 GPU.
4.2 Comparison
@@
Qualitative Comparison. As illustrated in Fig. 4 and Fig. 5, by incorporating semantics-guided objectives, Humansplat achieves more detailed and higher fidelity results against GTA [30] and LGM [29]. Although TeCH [21] also generates high-quality images, its SDS optimization often results in overly saturated multi-face outcomes [116, 117]. Moreover, while TeCH delivers the highest PSNR on the Thuman2.0 dataset, the LPIPS metric in Tab. 1, confirms that our method provides more realistic results. As shown in Fig. 6, Humansplat outperforms HumanSGD on in-the-wild images from Adobe Stock, effectively predicting 3D Gaussian Splatting properties without per-instance optimization. Notably, the results on in-the-wild images further demonstrate our strong generalization capability and our model provides superior textures on both the front and invisible regions, highlighting the effectiveness of incorporating appearance and geometric priors.